OpenStudy (anonymous):

a number added to 2 is 4 less than the number?

OpenStudy (mertsj):

Call the number x and add it to 2

OpenStudy (mertsj):

is means =

OpenStudy (mertsj):

how do you make something 4 less?

OpenStudy (anonymous):

\[x+2=x-4 \rightarrow 0 = -6 \rightarrow D.N.E\]There is no solution here for the number x here. @cgerm

OpenStudy (mertsj):

That is correct.

OpenStudy (mertsj):

And guess what? Everyone is involved in this question except the asker!!

OpenStudy (anonymous):

because this questions rediculous

OpenStudy (anonymous):

i know lol

OpenStudy (mertsj):

Unless it is purely an exercise in translation from English to mathematics